### v1

# LARES: Latent Reasoning for Sequential Recommendation

Abstract.
Sequential recommender systems have become increasingly important in real-world applications that model user behavior sequences to predict their preferences. However, existing sequential recommendation methods predominantly rely on non-reasoning paradigms, which may limit the model’s computational capacity and result in suboptimal recommendation performance. To address these limitations, we present LARES, a novel and scalable LAtent REasoning framework for Sequential recommendation that enhances model’s representation capabilities through increasing the computation density of parameters by depth-recurrent latent reasoning. Our proposed approach employs a recurrent architecture that allows flexible expansion of reasoning depth without increasing parameter complexity, thereby effectively capturing dynamic and intricate user interest patterns. A key difference of LARES lies in refining all input tokens at each implicit reasoning step to improve the computation utilization. To fully unlock the model’s reasoning potential, we design a two-phase training strategy: (1) Self-supervised pre-training (SPT) with dual alignment objectives; (2) Reinforcement post-training (RPT). During the first phase, we introduce trajectory-level alignment and step-level alignment objectives, which enable the model to learn recommendation-oriented latent reasoning patterns without requiring supplementary annotated data. The subsequent phase utilizes reinforcement learning (RL) to harness the model’s exploratory ability, further refining its reasoning capabilities. Comprehensive experiments on real-world benchmarks demonstrate our framework’s superior performance. Notably, LARES exhibits seamless compatibility with existing advanced models, further improving their recommendation performance.
Sequential Recommendation, Latent Reasoning
Equal contribution.
✉ Corresponding author.
journalyear: 2018copyright: acmcopyrightconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYdoi: 0000000.0000000isbn: 978-1-4503-XXXX-X/18/06price: 15.00ccs: Information systems Recommendation systems
1. Introduction
In the era of information explosion, recommender systems have emerged as indispensable components across real-world applications ranging from e-commerce platforms to online streaming services (Singer et al., 2022; Zheng et al., 2024). Among them, current research has increasingly focused on the analysis of evolving user behaviors for capturing latent intentions and sequential patterns to predict future interactions, which is termed as sequential recommendation. Recent years have witnessed significant advancements in this area, with notable methods like SASRec (Kang and McAuley, 2018) and BERT4Rec (Sun et al., 2019), which adopt transformer-based architectures for improved user behavior modeling.
To cope with the intricate and dynamic nature of user behaviors, extensive research (Zhai et al., 2024; Xu et al., 2025b; Zhang et al., 2024) has been conducted to scale the computational capacity of sequential recommenders, thereby strengthening their representational power. Previous works (Zhang et al., 2024; Zhai et al., 2024; Yan et al., 2025) have primarily pursued this goal through parameter scaling. However, these efforts have failed to replicate the success achieved in large language models (LLMs). This discrepancy primarily stems from unique challenges in recommendation systems, including inherent data sparsity and quality limitations that impede effective model scaling (Zhang et al., 2024).
Recent advances in large reasoning models (DeepSeek-AI et al., 2025; Team et al., 2025; Chen et al., 2025a) have demonstrated that scaling test-time computation is another effective approach to increasing the utilization of computing power and can significantly enhance the reasoning capabilities of LLMs. This suggests a promising approach to improving model performance by increasing the computation density for each parameter. There are two primary paradigms for test-time scaling in LLMs: one is explicit reasoning (Guo et al., 2024; Shao et al., 2024; Yang et al., 2024, 2025), where models verbalize intermediate reasoning steps (i.e., chain-of-thoughts (CoTs)) by generating meaningful tokens before producing final answers, and the other is latent reasoning (Hao et al., 2024; Xu et al., 2025a; Geiping et al., 2025), where models perform multi-step implicit reasoning in the latent space without generating explicit tokens. While most LLMs adopt explicit CoT reasoning, this approach faces challenges in recommendation systems. Unlike LLMs operating in dense textual spaces, most sequential recommenders are confined to sparse item ID spaces. This fundamental difference makes it hard to define meaningful reasoning steps like CoTs and provide supervision signals for training (Lightman et al., 2024; Luo et al., 2024). Therefore, we adopt the latent reasoning approach to scaling sequential recommenders. Recent work like ReaRec (Tang et al., 2025) has demonstrated the potential of this paradigm through autoregressive generation of implicit reasoning tokens. Despite its effectiveness, this method remains computationally suboptimal as it only enriches one token per reasoning step.
To tackle this problem, we aim to fully unleash computation by leveraging all input tokens at each reasoning step. Our approach is inspired by the recently proposed depth-recurrent latent reasoning paradigm in LLMs (Geiping et al., 2025), which leverages a middle recurrent architecture to realize test-time scaling by iterating a middle recurrent block.
To this end, we propose LARES, a novel and scalable LAtent REasoning framework for Sequential recommendation that enables flexible test-time scaling by thinking in continuous latent spaces. Our approach adopts a recurrent architecture comprising two key components: a pre-block for initial processing and a core-block for iterative refinement. The core-block supports arbitrary iteration depth, allowing for dynamic computation scaling while improving computational density (i.e., the amount of computation per parameter). To fully unlock the model’s reasoning capabilities, we develop a two-stage training strategy, the self-supervised pre-training (SPT) and reinforcement post-training (RPT). During the SPT stage, we propose the trajectory-level alignment and step-level alignment objectives to equip the model with recommendation-oriented latent reasoning patterns. Specifically, in trajectory-level alignment, we want to achieve mutual enhancement between reasoning actions for the same target. In step-level alignment, we aim to improve the thinking coherence among different intermediate steps. During the RPT stage, we employ reinforcement learning to further refine the model’s reasoning capabilities for recommendation.
In summary, our work makes the following main contributions:
$\bullet$ We propose LARES, a novel scalable latent reasoning approach for sequential recommendation that leverage all the input tokens to perform multi-step reasoning in latent space with arbitrary depth.
$\bullet$ We design a two-stage training strategy including SPT and RPT to fully unleash the model’s reasoning capabilities. During SPT, we introduce trajectory-level and step-level alignment to improve reasoning coherence. For RPT, we leverage RL to further improve recommendation performance via task-aligned rewards.
$\bullet$ Extensive experiments on real-world benchmarks validate LARES’s superiority, demonstrating its effectiveness and seamless compatibility with existing sequential recommenders.
Figure 1. Overall framework of LARES. $\mathcal{A}$ and $\mathcal{C}$ denote the pre-block and core-block, repsectively. “TLA” and “SLA” represent the Trajectory-Level Alignment and Step-Level Alignment.
2. Methodology
In this section, we present our scalable latent reasoning approach for sequential recommendation, named LARES, which is illustrated in Figure 1.
2.1. Overview
Sequential Recommendation. In recommender systems, there are a set of users $\mathcal{U}$ and a set of items ${\mathcal{V}}$. Let $M=|\mathcal{U}|$ and $N=|\mathcal{V}|$ denote the size of the user set and item set. The behavior record for each user $u\in\mathcal{U}$ is defined as an interaction sequence $S_{u}=[v_{1},\dots,v_{n}]$ at the time step $n$, where items are arranged in chronological order. As a typical sequential recommendation setting, traditional sequential recommenders (Zhou et al., 2022; Kang and McAuley, 2018; Zhou et al., 2020) encode the interaction sequence by direct inference to obtain sequential behavior representations. The recommender then predicts the next item the user is most likely to interact with based on the similarities between the encoded user representations and candidate item representations. The objective of next item prediction can be formally written as:
(1) $\max_{\Theta}P(v_{n+1}|\{v_{1},\dots,v_{n}\},\Theta),$
where $\Theta$ denotes the parameters of the sequential recommender.
Latent Reasoning For Sequential Recommendation. Traditional sequential recommenders typically adopt a straightforward inference pattern to directly provide recommendations, which struggle with complex recommendation tasks. In contrast, we propose a latent reasoning sequential recommender which is capable of thinking in a continuous space before making recommendations. This multi-step thinking paradigm allows the model to iteratively refine its inference in latent space and obtain more accurate user interests, simulating human-like mental thinking processes when addressing challenging problems. Specifically, the model progressively derives a series of intermediate thoughts, denoted as $T_{u}=[t_{1},...,t_{k}]$ where $k$ is the number of total reasoning steps, and makes final recommendations conditioned on these latent thoughts as well as the behavior sequence. Formally, the objective of latent reasoning sequential recommenders can be formulated as:
(2) $\max_{\Theta}P(v_{n+1},T_{u}|S_{u},\Theta)=P(v_{n+1}|T_{u},S_{u},\Theta)\cdot P(T_{u}|S_{u},\Theta).$
Scaling Latent Reasoning With Recurrent Depth. We present an overview of our proposed framework LARES, as depicted in Figure 1. The architecture adopts a depth-recurrent design consisting of a pre-block for mapping initial features into latent space and a core-block iterable to arbitrary depths, enabling flexible test-time scaling without additional parameters. To facilitate effective latent reasoning, we adopt a two-stage training paradigm for LARES that progressively learns how to think and model genuine user interests in the latent space. This approach draws inspiration from established practices in natural language processing, where pre-training enables LLMs to acquire fundamental knowledge, while reinforcement learning in the post-training phase enhances their reasoning capacity for complex tasks like mathematics, coding and scientific problems, even without supervised fine-tuning (SFT) data (Jaech et al., 2024; DeepSeek-AI et al., 2025; Team et al., 2025; Chen et al., 2025a). While some studies (Dang et al., 2025; Yue et al., 2025) suggest that reinforcement learning may not expand LLMs’ capability boundaries but instead accelerate solution discovery, these findings collectively emphasize the foundational importance of base model ability. Accordingly, our framework first employs self-supervised pre-training to equip LARES with core user interest inference capabilities through iterative latent reasoning, then applies reinforcement post-training to further enhance its reasoning performance by exploring diverse latent thought patterns.
The remainder of this paper is organized as follows: In Section 2.2, we introduce the depth-recurrent architecture of the latent reasoning recommender. Section 2.3 presents the details of self-supervised pre-training, including trajectory-level alignment and step-level alignment, enabling the model to possess the basic ability to infer user interests through a multi-step latent thinking process. In Section 2.4, we describe the process of reinforcement post-training, which leverages reinforcement learning algorithms to encourage the model to explore more thinking paths and further enhance the recommendation performance.
2.2. Latent Reasoning with A Recurrent Architecture
Our proposed framework, LARES, is a depth-recurrent sequential model consisting of multiple transformer layers, which can be adapted to other advanced sequential recommendation architectures,e.g., FMLPRec (Zhou et al., 2022), BSARec (Shin et al., 2024), and TedRec (Xu et al., 2024), as proved in the experiments in Section 3.4.1. These layers are primarily divided into two groups, the pre-block $\mathcal{A}$ and the core-block $\mathcal{C}$. The pre-block transforms the item embeddings into the latent space and then the core-block performs a multi-step reasoning on the latent representations in a recurrent pattern. Finally, the last latent thought representation is employed for subsequent similarity computation. The most significant difference of LARES lies in the recurrent design of the core block, allowing the model to dive into a deeper thinking of user behaviors without introducing extra parameter burdens. This design is inspired by the findings that the reasoning performance is up to a large effective depth but not necessarily many parameters in natural language processing (Saunshi et al., 2025; Geiping et al., 2025). The advantages of the recurrent architecture are twofold: on the one hand, it enhances the model’s computational expressiveness with no extra parameter memory cost; on the other hand, it enables more flexible inference scaling by controlling the recurrent depth.
Suppose the historical behavior sequence of user $u$ is $S_{u}=[v_{1},\dots,v_{n}]$. Given the item embedding table $\bm{E}\in\mathbb{R}^{N\times d}$, the sequence $S_{u}$ is first transformed into item embeddings $\bm{E}_{u}=[\bm{e}_{1};\dots;\bm{e}_{n}]$ by table lookup, where $d$ is the embedding dimension and [;][;][ ; ] denotes the concatenation operation. First, $\bm{E}_{u}$ along with the position embeddings $\bm{P}\in\mathbb{R}^{n\times d}$ is fed into the pre-block $\mathcal{A}$ to produce the initial latent item representation $\bm{H}\in\mathbb{R}^{t\times d}$, which is written as:
(3) $\bm{H}=\mathcal{A}\left(\bm{E}_{u}+\bm{P}\right).$
Second, the recurrent core-block $C$ performs iterative latent reasoning taking $\bm{H}$ and the latent thought of the previous step $\bm{T}_{i}$ as inputs, formulated as:
(4) $\bm{T}_{i}$ $=\mathcal{C}\left(\operatorname{LN}\left(f\left(\bm{T}_{i-1},\bm{H}\right)\right)\right),\ \ \text{for}\ i\in\{1,\dots,k\},$
(5) $\bm{T}_{0}$ $=\omega*\bm{z},\ \bm{z}\sim\mathcal{N}(0,\sigma_{1}^{2}I),$
where $i$ and $k$ denote the $i$th step and the number of total reasoning steps, $f$ is the aggregation function e.g., addition and concatenation, $\operatorname{LN}[\cdot]$ denotes the layer normalization operation and $\sigma_{1}$ is the standard deviation of the normal distribution for initializing the random state $\bm{T}_{0}$ and $\omega$ is a rescale coefficient. The input $\bm{H}$ to core-block $\mathcal{C}$ in every reasoning step functions as a residual connection to ensure stable gradient backpropagation. Finally, the final latent thought state corresponding to the last item $v_{n+1}$ is regarded as the final user preference representation $\bm{p}\in\mathbb{R}^{d}$ denoted as $\bm{p}=\bm{T}_{k}[-1]$. The output $\bm{p}$ is then used to compute the recommended probabilities of candidate items $\hat{y}_{i}$, and we adopt the widely used cross-entropy loss as the recommendation objective, which is formulated as:
(6) $\mathcal{L}_{\operatorname{Rec}}=-\log\hat{y}_{n+1}=-\log\frac{\exp({\bm{p}\cdot\bm{e}^{T}_{n+1})}}{\sum_{i=1}^{N}\exp({\bm{p}\cdot\bm{e}^{T}_{i}})},$
where $\bm{e}_{n+1}$ denotes the target item embedding of $v_{n+1}$ that user $u$ will interact with at the time step $n+1$.
2.3. Self-Supervised Pre-training for Thinking Adaptation
In the self-supervised pre-training (SPT) stage, LARES learns to perform recommendation-oriented latent reasoning for user interest modeling. However, relying solely on the recommendation objective $\mathcal{L}_{\text{Rec}}$ is not enough to ensure effective training because it cannot provide sufficient signals for the intermediate latent thinking process of LARES. To alleviate this problem, we propose two self-supervised optimization objectives, i.e., trajectory-level alignment and step-level alignment, to provide auxiliary supervision.
To ensure flexible scaling of inference-time reasoning, we sample the iteration count $k$ per training step from a log-normal Poisson distribution. For a target mean iteration count $\bar{k}+1$ (where $\bar{k}\in\mathbb{N}$) and variance $\sigma_{2}$, the sampling procedure is defined as follows:
(7) $\xi$ $\sim\mathcal{N}\left(\log(\bar{k})-\frac{1}{2}\sigma_{2}^{2},\sigma_{2}^{2}\right),$
(8) $k$ $\sim\mathcal{P}(e^{\xi})+1,$
where $\mathcal{N}$ and $\mathcal{P}$ denote the normal and Poisson distributions, respectively. In our experiments, we set $\sigma_{2}=0.5$. This distribution predominantly samples values below $\bar{k}$, while retaining a heavy tail that occasionally yields significantly higher iteration counts. To mitigate excessive memory consumption, we truncate $k$ such that $k=\min(k,3\bar{k})$.
2.3.1. Trajectory-Level Alignment
To further strengthen LARES’s reasoning capability, we introduce trajectory-level alignment to leverage complementary strengths from different reasoning trajectories. During inference, stochastic elements like random initialization and dropout naturally lead to varied reasoning paths across forward passes. We capitalize on this behavior by encouraging knowledge transfer between high-quality reasoning processes. Specifically, we align the final outputs of independent reasoning trajectories for identical input sequences. Following the approach in DuoRec (Qiu et al., 2022), we expand positive instance diversity by treating training sequences sharing the same target item as auxiliary positive views. However, our experiments reveal that aligning two reasoning outcomes with different step lengths adversely impacts model performance. We posit that this occurs because the inconsistency in reasoning steps causes a misalignment between long-chain and short-chain reasoning. Specifically, forcing short-chain reasoning to capture the richer and more complex patterns in longer reasoning processes is inherently challenging. As a result, the model tends to degenerate toward short-chain reasoning, ultimately compromising its effectiveness. To address this, we ensure consistent reasoning steps within each positive pair. Formally, given two positive sequences $S_{u}$ and $\hat{S}_{u}$ and a shared reasoning step $k$, LARES produces final preference representations $\bm{p}=\bm{T}_{k}[-1]$ and $\hat{\bm{p}}=\hat{\bm{T}}_{k}[-1]$. We achieve the alignment between them based on the InfoNCE loss. The trajectory-level alignment objective is formulated as:
(9) $F(\bm{x},\bm{y}^{+},\mathcal{B}_{\bm{y}})$ $=-\log\frac{\exp(s(\bm{x},\bm{y}^{+})/\tau)}{\sum_{\bm{y}\in\mathcal{B}_{y}}\exp(s(\bm{x},\bm{y})/\tau)},$
(10) $\mathcal{L}_{\text{TLA}}$ $=\frac{1}{2}\left(F(\bm{p},\hat{\bm{p}},\mathcal{R}_{\hat{\bm{p}}})+F(\hat{\bm{p}},\bm{p},\mathcal{R}_{\bm{p}})\right),$
where $F(\cdot,\cdot,\cdot)$ denotes InfoNCE loss function, $\mathcal{R}_{\bm{p}}$ and $\mathcal{R}_{\hat{\bm{p}}}$ represent sample sets containing both positive and negative instances.
2.3.2. Step-Level Alignment
Given a historical interaction sequence $S_{u}$ and a sampled iteration number $k$, LARES generates a sequence of latent thought representations $\bm{T}_{u}=[\bm{t}_{1},\dots,\bm{t}_{k}]$. Ideally, these latent thoughts are progressively refined to converge toward the true user preference distribution. While the model is expected to produce increasingly accurate latent representations as reasoning advances, intermediate states may occasionally diverge from the desired trajectory, leading to suboptimal or counterproductive reasoning steps. To address this issue, we introduce step-level alignment to enforce coherence between intermediate states and the final output. Specifically, we uniformly sample an intermediate step $b$ from $\{1,\dots,k-1\}$ and optimize the alignment between $\bm{o}_{b}$ and $\bm{o}_{k}$ using an InfoNCE loss:
(11) $b$ $\sim\operatorname{Unif}\{1,2,\dots,k-1\},$
(12) $\mathcal{L}_{\text{SLA}}$ $=\frac{1}{2}\left(F(\bm{t}_{b},\bm{t}_{k},\mathcal{B}_{k})+F(\bm{t}_{k},\bm{t}_{b},\mathcal{B}_{b})\right),$
where $\operatorname{Unif}$ denotes the uniform distribution, $s(\cdot,\cdot)$ is a similarity metric (e.g., cosine or dot product), $\mathcal{B}_{i},i\in\{b,k\}$ contains the $i$th step latent reasoning representations of all instances in the same batch, and $\tau$ is a temperature coefficient controlling the distribution sharpness.
The overall objective for self-supervised pre-training is written as:
(13) $\mathcal{L}_{\text{SPT}}=\mathcal{L}_{\text{Rec}}+\alpha\mathcal{L}_{\text{TLA}}+\gamma\mathcal{L}_{\text{SLA}},$
where $\alpha$ and $\gamma$ denote hyper-parameters to balance the weights among different objectives during optimizations, respectively.
2.4. Reinforcement Post-training for Thinking Exploration
During the self-supervised pre-training stage, the model acquires latent reasoning patterns for sequential recommendation tasks. However, this stage suffers from limited thinking exploration due to the absence of supervisory signals that guide the model in distinguishing between ‘good’ and ‘not good’ reasoning steps. Consequently, the model’s exploratory potential remains underutilized. To mitigate this limitation, we introduce a reinforcement learning-based post-training approach that enhances the model’s reasoning capabilities through learning from experiences of high-quality reasoning trajectories. The subsequent sections introduce the reinforcement learning algorithm, reward design, and data selection startegy.
2.4.1. Reinforcement Learning Algorithm
To balance performance and computational cost, we employ the Group Relative Policy Optimization (GRPO) algorithm (Shao et al., 2024) during reinforcement post-training. For each input (i.e., user interaction sequences $S_{u}$ in our case), GRPO samples a group of $G$ rollouts $\{y_{1},\dots,y_{G}\}$ from the old policy $\pi_{\theta_{\text{old}}}$ where $G$ denotes the group size. The current policy $\pi_{\theta}$ (i.e., the base model) is then updated by maximizing a surrogate reward, regularized by the KL divergence from a reference policy $\pi_{\theta_{\text{ref}}}$ (i.e., the initial pre-trained model). The formulation of GRPO in our scenario is partially different from the original setting. Specifically, the output $y_{i}$ differs from conventional LLM outputs (i.e., response tokens). Instead, it represents recommended items conditioned on latent thoughts at different steps, denoted as $t^{i}=\{t_{j}^{i}\}_{j=1}^{k}$. To align with sequential recommendation objectives, we assume the target item (rather than the highest-probability candidate) is recommended at each step, thereby explicitly maximizing the target item likelihood. The rollout probability is accordingly defined as the joint probabilities of recommending target items for all latent thinking states: $\pi(y_{i}|S_{u})=\prod_{j=1}^{k}\pi(v_{n}|S_{u},t^{i}_{j})$ where $v_{n}$ denotes the target item. During training, we perform one-step updates on rollouts, maintaining $\pi_{\theta}=\pi_{\theta_{\text{old}}}$. Overall, the GRPO objective is formally expressed as:
(14) $\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{S_{u}\sim D,\{t^{i}\}_{i=1}^{G}\sim\pi_{\theta}}\left[\frac{1}{G}\sum_{i=1}^{G}\left(A_{i}-\beta\ \mathbb{D}_{\text{KL}}\left(\pi_{\theta}||\pi_{{\text{ref}}}\right)\right)\right],$
(15) $\mathbb{D}_{\text{KL}}\left(\pi_{\theta}||\pi_{{\text{ref}}}\right)=\prod_{j=1}^{k}\frac{\pi_{{\text{ref}}}(v_{n}|S_{u},t^{i}_{j})}{\pi_{\theta}(v_{n}|S_{u},t^{i}_{j})}-\sum_{j=1}^{k}\log\frac{\pi_{{\text{ref}}}(v_{n}|S_{u},t^{i}_{j})}{\pi_{\theta}(v_{n}|S_{u},t^{i}_{j})}-1.$
2.4.2. Reward Design
To maintain strict alignment with recommendation objectives, we directly employ standard recommendation metrics (namely NDCG@K and Recall@K) as reward signals. The reward of each rollout is determined by the recommendation results of the last reasoning step and shared by all previous latent reasoning steps. For $G$ rollout trajectories $\{t^{1},\dots,t^{G}\}$, we first obtain the last latent thought states corresponding to the last item as the final user representations $\{\bm{p}_{1},\dots,\bm{p}_{G}\}$ and obtain the recommendation probability distribution for each rollout by computing the similarity between $\bm{p}_{i}$ and item embedding table $\bm{E}$, denoted as $P_{i}=\operatorname{softmax}(\bm{p}_{i}\bm{E}^{T})\in\mathbb{R}^{N},i\in\{1,\dots,G\}$. Then we can obtain the item ranking list based on the probability distribution for every rollout, denoted as $\operatorname{IL}_{i}=\operatorname{argsort}P_{i}$. The reward function is formally written as:
(16) $r_{i}=m(v_{n},\operatorname{IL}_{i}),\quad m\in\{\text{NDCG}@K,\text{Recall}@K\}.$
The advantage value $A_{i}$ is computed as the z-score normalized reward within each group:
(17) $A_{i}=\frac{r_{i}-\operatorname{mean}(\{r_{1},\dots,r_{G}\})}{\operatorname{std}(\{r_{1},\dots,r_{G}\})}.$
2.4.3. Data Selection
To improve the training efficiency of Reinforcement Post-Training (RPT), we propose a data selection strategy that filters out hard training samples to reduce the size of training datasets. Specifically, we exclude instances where the pre-trained model fails to rank the target item within the top-100 candidates across three independent inference trials. This selection criterion is empirically justified, as we observe that such challenging instances contribute minimal reward signals during training and consequently have negligible impact on final model performance.
2.5. Time Complexity
As described in Section 2.2, the LARES framework comprises two key components: a pre-block and a core-block. To illustrate the computational complexity, we adopt the transformer architecture as a representative example due to LARES’s compatibility with diverse model designs. The primary computational overhead in each transformer layer stems from the Multi-Head Self-Attention (MHSA) and Feed-Forward Network (FFN) modules, with complexities of $\mathcal{O}(N^{2}d+Nd^{2})$ and $\mathcal{O}(Nd^{2})$ respectively, where $N$ represents the sequence length and $d$ denotes the hidden dimension. Let $L_{1}$ and $L_{2}$ denote the number of transformer layers in the pre-block and core-block, respectively. Consequently, their computational complexities can be expressed as $\mathcal{O}(L_{1}(N^{2}d+Nd^{2}))$ and $\mathcal{O}(L_{2}(N^{2}d+Nd^{2}))$. For a reasoning process involving $K$ iterative steps, the overall time complexity of LARES scales as $\mathcal{O}((L_{1}+L_{2}K)(N^{2}d+Nd^{2}))$.
3. Experiments
In this section, we conduct extensive experiments and analysis to empirically demonstrate the effectiveness of LARES.
3.1. Experiment Setup
3.1.1. Dataset
We assess our proposed method on four subsets derived from the latest Amazon 2023 review dataset: “Musical Instruments”, “Video Games”, “Baby Products”, and “Industrial & Scientific”. Following prior studies (Zhou et al., 2020, 2022; Xu et al., 2024), we employ 5-core filtering to exclude inactive users and unpopular items with fewer than five interactions to ensure a robust evaluation. All user interactions are grouped by user ID and sorted chronologically. We truncate behavior sequences to a maximum of 20 items per user. Table 1 summarizes the key statistics of the preprocessed datasets.
3.1.2. Baseline Models
To conduct a comprehensive evaluation of LARES’s performance, we compare it with multiple sequential recommendation (SR) baselines, including both reasoning-based and non-reasoning approaches:
(1) Non-Reasoning SR methods:
  • GRU4Rec (Hidasi et al., 2016a) employs GRUs to capture user behavior patterns.
  • SASRec (Kang and McAuley, 2018) is a transformer-based model utilizing unidirectional multi-head self-attention to encode user interaction sequences.
  • BERT4Rec (Sun et al., 2019) is a bidirectional self-attentive model that employs masked prediction for sequence modeling.
  • FMLP-Rec (Zhou et al., 2022) replaces traditional self-attention with filter-enhanced MLPs to improve behavior modeling.
  • BSARec (Shin et al., 2024) leverages Fourier transforms to capture both high- and low-frequency information in user behavior sequences.
  • CL4SRec (Xie et al., 2022) first introduces contrastive learning for sequential recommendation through three augmentation strategies: item masking, reordering, and cropping.
  • DuoRec (Qiu et al., 2022) combines unsupervised model-level dropout augmentation with hard positive sample selection.
(2) Reasoning-based SR methods:
  • ERL (Tang et al., 2025) enhances sequential recommendations by aggregating multi-step implicit reasoning states.
  • PRL (Tang et al., 2025) improves reasoning capabilities for sequential recommendation through contrastive learning with noise-disturbed positive views and temperature annealing.
  • PRL++ extends PRL by incorporating DuoRec’s sampling strategy.
Table 1. Statistics of the preprocessed datasets. Avg.L represents the average length of user interaction sequences.
Dataset #Users #Items #Actions Avg.L Sparsity
Instrument 57,439 24,587 511,836 8.91 99.964%
Scientific 50,985 25,848 412,947 8.10 99.969%
Game 94,762 25,612 814,586 8.60 99.966%
Baby 150,777 36,013 1,241,083 8.23 99.977%
Table 2. Performance comparison of different methods. The best and second-best results are indicated in bold and underlined font, respectively. “*” denotes that the improvements are statistically significant with $p<0.01$ in a paired t-test setting.
Dataset Metric GRU4Rec BERT4Rec SASRec FMLP-Rec CL4SRec DuoRec BSARec ERL PRL PRL++ LARES
[…]
3.1.3. Evaluation Settings
We evaluate the sequential recommendation task using two standard metrics: Recall@$K$ and NDCG@$K$, with $K\in\{5,10,20\}$. Following prior work (Zhou et al., 2020; Rajput et al., 2023; Liu et al., 2025), we adopt a leave-one-out strategy. Specifically, for each user interaction sequence, we use the most recent interaction as the test instance, the second-last interaction for validation, and all remaining historical interactions for training. To ensure rigorous evaluation and mitigate potential biases from negative sampling, we perform full ranking over the entire item pool. The final results report the average metric scores across all test instances.
Table 3. Ablation studies of LARES on three datasets by selectively removing the Trajectory-Level Alignment (TLA), Step-Level Alignment (SLA), and Reinforcement Post-Training (RPT). “N@K” and “R@K” denote “NDCG@K” and “Recall@K”, respectively.
Variants Instrument Scientific Game
TLA SLA RPT R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10
0.0356 0.0571 0.0221 0.0290 0.0245 0.0395 0.0148 0.0196 0.0559 0.0903 0.0337 0.0448
[…]
3.1.4. Implementation
We implement LARES and all baseline models using PyTorch. To ensure a fair comparison, the batch size, embedding size and hidden size are set to 1024, 64 and 256, respectively. All models are optimized using the AdamW optimizer with a learning rate of 0.001. For self-attentive models, the number of attention heads is fixed at 2. For LARES, the number of layers for both the pre-block and core-block is selected from $\{1,2\}$, while the mean reasoning step $\bar{k}$ is chosen from $\{3,4,6\}$. The hyperparameters for latent reasoning state initialization $\sigma_{1}$ and $\omega$ are set to 1 and 3, respectively. In SPT, we use a learning rate of 0.001 and a dropout rate of 0.5. The coefficients for trajectory-level alignment $\alpha$ and step-level alignment $\gamma$ are tuned within $\{0.1,0.2,0.3\}$ and $\{0.1,0.3,0.5,0.7\}$, respectively. In RPT, the learning rate is selected from $\{0.0005,0.0003,0.0001\}$, and $\beta$ is tuned from $\{0.5,1.0\}$. The rollout number $G$ is fixed at 4. The reward function is selected from $\{\operatorname{Recall@5},\operatorname{Recall@10}\}$. The non-reasoning SR baselines are implemented using RecBole (Zhao et al., 2021, 2022), an open-source recommendation library, while reasoning-based SR models are reproduced from their official source codes. To mitigate overfitting, we employ early stopping, terminating training if NDCG@10 on the validation set shows no improvement for 10 consecutive epochs.
3.2. Overall Performance
We evaluate the performance of our proposed LARES framework by comparing it with various baseline methods across four real-world benchmark datasets. The comprehensive experimental results are presented in Table 2, from which we draw the following key observations:
  • For non-reasoning models, SASRec surpasses BERT4Rec and GRU4Rec across all datasets, confirming the superiority of transformer architectures in modeling user behavior sequences. And BERT4Rec underperforms both GRU4Rec and SASRec, suggesting that bidirectional self-attention mechanisms incorporating future context may be suboptimal for sequential recommendation tasks. The filter-enhanced models (i.e., FMLP-Rec and BSARec) perform better than SASRec, demonstrating the effectiveness of noise reduction in the frequency domain for capturing sequential patterns. Contrastive learning approaches (i.e., CL4SRec and DuoRec) consistently outperform traditional ID-based baselines, showing that contrastive regularization can enhance the SR model’s representation learning. DuoRec’s superior performance over CL4SRec stems from its robust model-level dropout augmentation strategy compared to sequence-level augmentations.
  • For reasoning models, ERL achieves superior performance than SASRec across most datasets, indicating that performing multi-step reasoning can better capture the user preference. PRL consistently outperforms ERL demonstrating that noise-disturbed reasoning outcomes as contrastive signals can enhance the model’s ability to extract critical sequence information. PRL++ shows significant improvements over PRL, proving that incorporating hard positive samples effectively strengthens the effectiveness of contrastive learning.
  • Our proposed framework LARES outperforms all baselines, including both non-reasoning and reasoning models, by a large margin across all evaluation metrics on four datasets. This consistent superiority underscores its advanced reasoning capabilities for sequential recommendation tasks. LARES enables multi-step latent reasoning through core-block iteration at a arbitrary depth, allowing flexible computation scaling by increasing the computational density of parameters. To fully exploit its latent reasoning potential, we design two training stages, self-supervised pre-training to instill latent reasoning patterns tailored for sequential recommendation tasks and reinforcement post-training to further stimulate its reasoning capabilities by encouraging thinking exploration. These results also validate the effectiveness of recurrent-depth latent reasoning for sequential recommendation tasks.
3.3. Ablation Studies
We conduct an ablation study to evaluate the contribution of each key component in LARES: Trajectory-Level Alignment (TLA), Step-Level Alignment (SLA), and Reinforcement Post-Training (RPT). As shown in Table 3, we gradually remove these components and have the following findings:
  • The removal of any single component (TLA, SLA, or RPT) consistently degrades performance across all datasets. The complete removal of all components (leaving only the backbone model with latent reasoning) yields the worst performance, demonstrating that both alignment objectives in the pre-training stage are essential and RPT provides additional improvements.
  • The variant incorporating both TLA and SLA outperforms the version with TLA alone, confirming that these alignment strategies provide complementary benefits: TLA enhances trajectory-level reasoning consistency and SLA improves step-level reasoning coherence.
  • The relative importance of TLA and SLA varies across domains. Specifically, TLA shows greater impact on Instrument and Game datasets, suggesting hard positive samples are particularly informative in these domains. SLA proves more effective for Scientific dataset, indicating that reasoning coherence is especially valuable in Scientific scenario.
Figure 2. Performance comparison of LARES across different backbone architectures on Instrument and Scientific. ‘Base’ indicates the original backbone model. ‘LR+SPT’ refers to the backbone enhanced with our proposed latent reasoning module and pre-training. ‘LARES’ denotes the full implementation incorporating all proposed components.
3.4. Further Analysis
3.4.1. Compatibility with Different Backbones
To evaluate the compatibility of LARES with different architectures, we conduct experiments using three advanced backbones in addition to SASRec (namely FMLP-Rec, BSARec, and TedRec (Xu et al., 2024)). The first two are conventional ID-based sequential recommendation models, whereas TedRec is a text-enhanced model that incorporates both ID embeddings and text embeddings derived from item metadata using a pretrained language model (i.e., Sentence-T5 in our case). Figure 2 presents the performance comparison of three model variants: (1) the backbone (Base), (2) the backbone enhanced with latent reasoning through pre-training (LR+SPT), and (3) the complete framework (LARES). The experimental results demonstrate that LARES consistently improves all four backbones, achieving performance gains of 15.3% and 21.4% in NDCG@10 for FMLP-Rec on the Instrument and Scientific datasets, respectively. For BSARec, LARES yields over 10% improvement across all metrics on both datasets. Notably, even the state-of-the-art TedRec model benefits from our framework, showing average improvements of 10.9% and 7.9% on the respective datasets. These results substantiate the great compatibility of our latent reasoning paradigm. Furthermore, we observe a consistent performance ranking across all metrics and datasets: Base ¡ LR+SPT ¡ LARES. This hierarchy validates the effectiveness of our proposed two-stage training strategy.
Figure 3. Performance of different reasoning steps on Instrument and Scientific. ‘SPT’ denotes the latent reasoning model with only pre-training. ‘LARES’ denotes the model with both pre-training and post-training.
Figure 4. Performance of different alignment coefficients $\alpha$ and $\gamma$ on Instrument and Scientific.
3.4.2. Influence of Reasoning Steps on Recommendation Performance
To analyze the influence of reasoning step, we vary the number of reasoning steps $\bar{k}$ in $\{1,2,3,4,6,8\}$ and evaluate the model on the Instrument and Scientific datasets. The results in Figure 3 reveal a consistent trend: performance initially improves with more reasoning steps but declines after reaching an optimal point, which is in line with the findings in ReaRec. This pattern suggests that moderate reasoning step enhances model performance by strengthening representation power through additional computation. However, excessively large $\bar{k}$ leads to performance degradation, likely because simple user interaction sequences do not require intensive reasoning-a phenomenon analogous to “overthinking” in NLP . The optimal reasoning steps are 3 for Instrument and 2 for Scientific, highlighting the importance of selecting an appropriate reasoning depth for optimal performance.
3.4.3. Influence of Alignment Coefficients
We examine the effects of trajectory-level alignment and step-level alignment coefficients by evaluating model performance on Instrument and Scientific datasets across varying values $\alpha\in$0.1, 0.2, 0.3, 0.4} and $\gamma\in$0.1, 0.3, 0.5, 0.7}. From the results in Figure 4, we can observe that model performance on both datasets fluctuates with increasing $\alpha$ values, peaking at 0.1 and reaching the lowest point at 0.4. This suggests that excessive trajectory-level alignment may interfere with sequential pattern learning. For $\gamma$, the performance on Instrument initially improves and then declines at higher $\gamma$ values. The performance on Scientific exhibits generally consistent improvement with increasing $\gamma$ values. These findings highlight the critical role of intra-step reasoning coherence in maintaining reasoning quality, while revealing differential sensitivity to alignment coefficients across datasets.
3.4.4. Influence of Reasoning Steps on Inference Time
To systematically assess the impact of reasoning steps on inference efficiency, we conduct a comparative analysis between LARES with varying reasoning steps and SASRec under equivalent computational budgets (FLOPs) on the Instrument dataset. LARES uses 2-layer pre-blocks and core-blocks, and the batch size is set to 1024 for both. As presented in Table 4 (where item embeddings are excluded from parameter calculations), LARES achieves higher computational complexity while maintaining identical parameter sizes through increased reasoning steps. For instance, with 4 reasoning steps, LARES attains an effective depth of 2+2×4 layers, comparable to a 10-layer SASRec architecture. The additional inference latency introduced by LARES remains moderate—approximately 10% higher than SASRec at comparable FLOPs. This computational overhead is acceptable in consideration of LARES’s substantial performance gains, which improve SASRec’s baseline by an average of 20% on Instrument as shown in Figure 2. Our findings demonstrate that LARES offers an advantageous trade-off between computational cost and model performance, suggesting strong potential for practical deployment.
Table 4. Comparison of inference time between LARES with different reasoning steps and SASRec with the same FLOPs on Instrument. The experiments are conducted on a single RTX3090 GPU.
Method Step Params FLOPs Inference Time
SASRec 1 4x 4x 1.10s
1 6x 6x 1.23s
1 8x 8x 1.40s
1 10x 10x 1.55s
LARES 1 2x+2x 4x 1.19s
2 2x+2x 6x 1.39s
3 2x+2x 8x 1.58s
4 2x+2x 10x 1.74s
4. Related Work
4.1. Sequential Recommendation
Sequential recommendation has become a prominent research area in recommender systems, with the objective of modeling latent patterns in user behavior sequences to predict the next item of interest. Early approaches (Rendle et al., 2010; He and McAuley, 2016) modeled user behaviors as Markov chains, focusing exclusively on item transition patterns. The advent of deep learning revolutionized this field, leading to the adoption of various neural architectures. These include convolutional neural networks (CNNs) (Tang and Wang, 2018; Yan et al., 2019; Yuan et al., 2019), recurrent neural networks (RNNs) (Hidasi et al., 2016a; Tan et al., 2016; Hidasi et al., 2016b; Quadrana et al., 2017), and graph neural networks (GNNs) (Chang et al., 2021; Wu et al., 2019). Recently, Transformer-based models (Kang and McAuley, 2018; Sun et al., 2019; Hao et al., 2023) have demonstrated superior performance in sequential behavior modeling. Several recent studies (Chen et al., 2022; Wang et al., 2023; Zhou et al., 2023; Liu et al., 2024) have proposed to enhance Transformer architectures. FMLP-Rec (Zhou et al., 2022) replaces self-attention with filter-enhanced MLPs to reduce noise in user preference modeling. However, these ID-based methods often suffer from cold-start problems. To address this, alternative approaches incorporate item textual metadata to enrich representations (Hou et al., 2022; Xu et al., 2024; Liu et al., 2025). UniSRec employs multi-domain textual data with MoE-enhanced adapters to learn universal sequential representations. TedRec achieves sequence-level fusion of textual and ID representations through contextual convolution. Despite these advancements, the test-time scaling in sequential recommendation is underexplored. In this paper, we propose a new latent reasoning paradigm for sequential recommendation that leverages all input tokens to perform multi-step reasoning in latent space with arbitrary depth.
4.2. Reasoning Models
Recent advances in test-time scaling (DeepSeek-AI et al., 2025; Chen et al., 2025a, b) have shifted the research focus in GenAI from large language models (LLMs) to large reasoning models (LRMs). LRMs excel in complex reasoning tasks through deep thinking capabilities, as demonstrated by the powerful reasoning systems like OpenAI-o1 and DeepSeek-R1. These models employ explicit long Chain-of-Thought mechanisms to generate extensive reasoning tokens before producing final answers (Yang et al., 2024; Besta et al., 2024; Yang et al., 2025). However, their reliance on explicit reasoning poses some critical challenges, including excessive memory demands from long context windows and limited expressive power due to discrete language space constraints. To address these limitations, recent studies (Hao et al., 2024; Chen et al., 2024; Geiping et al., 2025; Xu et al., 2025a) have introduced latent reasoning models that perform implicit reasoning in continuous latent spaces, achieving greater efficiency. In recommender systems, some efforts have adapted reasoning mechanisms for sequential recommendation models (Tang et al., 2025; Zhang et al., 2025). For instance, ReaRec autoregressively generates latent reasoning tokens to refine user representations. STREAM-Rec integrates slow-thinking paradigms with TIGER-style generative recommenders by producing annotated reasoning tokens before final recommendation semantic tokens. In contrast, we propose LARES, a novel recurrent-depth latent reasoning framework for sequential recommendation. Unlike prior work, LARES iteratively refines all input tokens at each reasoning step. Additionally, it is seamlessly compatible with existing sequential recommendation models, further enhancing their performance.
5. Conclusion
In this work, we propose LARES, a scalable and novel latent reasoning framework for sequential recommendations. A significant difference between LARES and previous latent reasoning methods like ReaRec lies in that our method leverages all input tokens during the latent reasoning process instead of generating only one reasoning token for higher computation utilization. Our approach adopts a depth-recurrent architecture consisting of a pre-block for mapping initial features into the latent space and an iterable core-block, enabling flexible test-time computation scaling without extra parameters. To fully exploit the potential of latent reasoning, we design a two-stage training pipeline including self-supervised pre-training (SPT) and reinforcement post-training (RPT). In the SPT stage, we introduce the trajectory-level alignment and step-level alignment to improve the model’s reasoning capabilities from two different aspects. Trajectory-level alignment aims to align the reasoning outcomes of the user interaction sequences with the same target for mutual enhancement, while the target of step-level alignment is to improve the intra-step reasoning coherence to avoid divergences of intermediate reasoning steps. In the RPT stage, we leverage reinforcement learning to further stimulate model’s reasoning abilities by encouraging it to explore diverse reasoning paths. To fully align with the downstream tasks, we directly adopt the standard metrics as rewards. Extensive experiments have demonstrated the superior performance of LARES and its seamless compatibility with existing advanced SR models.
Notably, there remain some promising directions for future research:
  • Adaptive Reasoning Depth. While our current approach employs a fixed reasoning step for all user interaction sequences, we observe that not all sequences necessitate deep reasoning. Excessive reasoning steps may incur unnecessary computational overhead and even degrade recommendation performance. This raises an open question: How to adaptively determine the reasoning depth for different user behavior sequences?
  • Process Reinforcement Learning. In our RPT stage, we only consider the outcome rewards which are coarse-grained for the intermediate reasoning states. Therefore, incorporating process reinforcement learning could provide more precise guidance and improve the overall reasoning quality.